Find sum of 9 terms in a
arithmetic sequence is 207 sum of first 20 terms
is 900 . Find fifth and twenty fifth term Find
the sum of 100 terms

Answer by stanbon(75887)

If the sum of 9 terms in an arithmetic sequence is 207 and the sum of first 20 terms is 900 ,
Find fifth and twenty-fifth term
--------
S(9) = (a(1)+(a(1)+8d)*(9/2) = 207
S(20) = (a(1)+(a(1)+19d)*(20/2) = 900
--------
2(a(1)) + 8d = (2/9)207 = 2*23 = 46
2(a(1)) + 19d = 90
------
Subtract and solve for "d"::
11d = 44
d = 4
----
Then, solve for a(1)::
2*a(1) + 8d = 46
2*a(1) + 32 = 46
a(1) = 7
------------------------
Ans: a(5) = 7 + 4d = 7 + 16 = 23
a(25) = 7 + 24d = 7 + 96 = 103
----------------------------
Find the sum of 100 terms.
s(100) = (7 + a(100))*(50)
a(100) = 7 + 99*4 = 403
S(100) = 410*50 = 20500
-------------
Cheers,
Stan H.
---------------